### v1

@@ 6 Conclusions
In this work we proposed a novel multi-view photometric stereo method. Unlike most MVPS methods our approach explicity leverages per-pixel intensity renderings rather than relying mainly on estimated normals. We believe such approach is required for truly applicable and robust MVPS as the estimated normals are likely to fail on complex materials or geometries. We clearly demonstrate the benefit of leveraging intensities on a synthetic and real DiLiGenT-MV benchmark and the applicability of our method on the minimal 6 lighst case.
Appendix 0.A Appendix
This document contains supplementary material. It is organised as follows: DiLiGenT-MV data anomalies in Section 0.A.1 and visualisation of shadows and indirect reflectance on synthetic data in Section 0.A.2.
0.A.1 DiLiGenT-MV data anomalies
@@
0.A.2 DiLiGenT-MV-Synthetic visualisation
This section contains additional visualisations for the relationship between shadows and indirect reflectance which is obtained by investigating the synthetic version of DiLiGenT-MV. This is shown in Figure 8 where it is shown that in regions with a high number of shadows (i.e. ambient occlusion $\tilde{s}$ ), indirect reflectance is higher.
Figure 8: Visualisation of the relationship between shadows and indirect illumination (self reflection) on the synthetic version of DiLiGenT-MV. First column shows the ambient occlusion estimate $\tilde{s}$ and second and third column show the mean and standard deviation (over all lights) of the attenuation compensated indirect reflectance $i_{r}/a$.

### v2

@@ 6 Conclusions
In this work we proposed a novel multi-view photometric stereo method. Unlike most MVPS methods our approach explicity leverages per-pixel intensity renderings rather than relying mainly on estimated normals. We believe such approach is required for truly applicable and robust MVPS as the estimated normals are likely to fail on complex materials or geometries. We clearly demonstrate the benefit of leveraging intensities on a synthetic and real DiLiGenT-MV benchmark and the applicability of our method on the minimal 6 lights case.
Computational efficiency has been beyond the scope of this project, however future work can target that aspect though intergration with the NERFACC [25] framework.
Appendix A Appendix
This document contains a brief discussion of DiLiGenT-MV data anomalies in Section B.
Appendix B DiLiGenT-MV data anomalies